# Sums of four fourth power of primes

### v2

@@ Abstract.
$\ell=p_{1}^{4}+p_{2}^{4}+p_{3}^{4}+p_{4}^{4},$
where $p_{1},p_{2},p_{3},p_{4}$ are primes.For such $\ell$, in this paper we will use circle method and sieves to prove that the proportion of $\ell$ in positive integers is at least $\frac{1}{27241.64}$ .
@@ 1. Introduction
For any positive integers $m$ and $k$, the Warning-Goldbach Problem is to discuss every positive number $\ell_{m,k}$ can be expressed as the sum of the mthm-th powers of $k$ primes $p_{1},p_{2},\cdots,p_{k}$, i.e.
@@
When $m=1,k=2$, this is the famous Goldbach’s Conjecture, the difficult problem, which has not been successfully solved so far. But when $m=1$, $k=3$, the case of three prime numbers has been proven, which is the famous theorem of three prime numbers[9]. Furthermore, for higher-order terms like $m=3,k=4$, some scholars have already obtained some results. Let $\mathcal{L}_{3,4}$ be the set of integers $\ell_{m,k}$ that can be written as the formula (1.1) when $m=3,k=4$ . In 1949, Roth[5] showed that
@@
Futhermore, from 2001 to 2003, Ren[2][3]improved Roth’s theorem to the extent that
@@
The authors are interested in this problem especially for more cases $m$ and $k$. In this paper we consider the case $m=4,k=4$ and try to say something about it. Let $\ell$ be a positive integer, suppose $\ell$ can be expressed as the sum of the four powers of four prime numbers, i.e.
(1.2) $\ell=p_{1}^{4}+p_{2}^{4}+p_{3}^{4}+p_{4}^{4},$
@@
Theorem. $\beta=\frac{1}{27241.64}$ is acceptable in (1.3).
Note. Compared to the sums of powers of four cubes of primes, the density of sums of four fourth power of primes is much smaller, but if we expand the expression to the sums of more than four prime numbers, the density may become even higher, which is a problem worthy of further research.
@@ 2. Prerequisite Knowledge
Let NN be a large integer, $\delta$ be small enough and
$U=\left(\frac{N}{32(1+\delta)}\right)^{1/4},\quad V=U^{5/6}.$
@@
(2.1) $J(n)=\int_{-U^{4}L^{B}}^{U^{4}L^{B}}\Phi(\lambda,U)\overline{\Psi(\lambda,U)}\Big|\Psi(\lambda,U)\Big|^{2}\Big|\Psi(\lambda,V)\Big|^{4}e(-n\lambda)\>d\lambda.$
@@
(2.3) $T_{d}(n,q)=\sum_{\overset{a=1}{(a,q)=1}}^{q}\frac{S(q,ad^{3})C^{3}(q,a)\overline{C^{4}(q,a)}}{q\varphi^{7}(q)}e\biggl(-\frac{an}{q}\biggr)$
@@
(2.4) $S(q,a)=\sum_{m=1}^{q}e\biggl(\frac{am^{4}}{q}\biggr),\quad C(q,a)=\sum_{(m,q)=1}^{q}e\biggl(\frac{am^{4}}{q}\biggr).$
@@ 3. Some Lemmas
$p_{2},p_{5},p_{6}\sim U,\quad p_{3},p_{4},p_{7},p_{8}\sim V.$
For $0<\xi<1/16$ and $D=N^{\xi}$, take $\mathfrak{S}_{d}(n)$ and $J(n)$ from the above (2.1), (2.2), define $E_{d}(n)$ as follows:
@@
$\ll\int_{U^{-4}L^{B}}^{\infty}|\Phi(\lambda,U)||\Psi(\lambda,U)|^{3}|\Psi(\lambda,V)|^{4}\>d\lambda$
$\ll U^{-12}V^{-12}L^{-7}\int_{U^{-4}L^{B}}^{\infty}\min(U^{4},|\lambda|^{-1})^{4}\min(V^{4},|\lambda|^{-1})^{4}\>d\lambda$
$\ll V^{4}L^{-3B}.$
@@
$\int_{-\infty}^{\infty}\Phi(\lambda,U)\overline{\Psi(\lambda,U)}\big|\Psi(\lambda,U)\big|^{2}\big|\Psi(\lambda,V)\big|^{4}e(-n\lambda)\>d\lambda$
$=$ $\frac{1}{4}\int_{\mathcal{D}}\frac{d\nu_{1}\cdots d\nu_{4}du_{1}\cdots du_{3}}{\nu_{1}^{3/4}\cdots\nu_{4}^{3/4}u_{1}^{3/4}\cdots u_{4}^{3/4}\log\nu_{1}\cdots\log\nu_{4}\log u_{1}\cdots\log u_{3}},$
@@
$\mathcal{D}=\left\{(v_{1},\ldots,v_{4},u_{1},\ldots,u_{3})\colon V^{4}\leqslant v_{1},\ldots,v_{4}\leqslant 16V^{4},\ U^{4}\leqslant u_{1},\ldots,u_{4}\leqslant 16U^{4}\right\}$
and $u_{4}=n+v_{1}+v_{2}-v_{3}-v_{4}+u_{1}+u_{2}-u_{3}.$
@@
(3.1) $\begin{split}J(n)=&\int_{-\infty}^{\infty}\Phi(\lambda,U)\overline{\Psi(\lambda,U)}|\Psi(\lambda,U)|^{2}|\Psi(\lambda,V)|^{4}e(-n\lambda)d\lambda\\ &+\>O(V^{4}L^{-3B})\\ =&\frac{1}{4}\int_{V^{4}}^{16V^{4}}\>\frac{dv_{1}}{v_{1}^{3/4}\log v_{1}}\cdots\int_{V^{4}}^{16V^{4}}\>\frac{dv_{4}}{v_{4}^{3/4}\log v_{4}}\int_{U^{4}}^{16U^{4}}\>\frac{du_{1}}{u_{1}^{3/4}\log u_{1}}\\ &\times\int_{U^{4}}^{16U^{4}}\frac{du_{2}}{u_{2}^{3/4}\log u_{2}}\int_{\max(U^{4},x-16U^{4})}^{\min(16U^{4},x-U^{4})}\frac{du_{3}}{u_{3}^{3/4}(x-u_{3})^{3/4}\log u_{3}}\\ &+\>O(V^{4}L^{-3B}).\end{split}$
Let the last integral in the equation be denoted as II, where $x=n+v_{1}+v_{2}-v_{3}-v_{4}+u_{1}+u_{2}$, now we calculate the upper bound of II. Firstly, regarding $x\leqslant 2U^{4}$ or $x\geqslant 32U^{4}$, in both cases, the integral region does not exist, so we have $2U^{4}<x<32U^{4}$. In this case, we record as $u_{3}=xu$, then there is

### v5

# Sums of eight fourth power of primes

@@ Abstract.
$\ell=p_{1}^{4}+p_{2}^{4}+p_{3}^{4}+\cdots+p_{8}^{4},$
where $p_{1},p_{2},p_{3},\cdots,p_{8}$ are primes.For such $\ell$, in this paper we will use circle method and sieves to prove that the proportion of $\ell$ in positive integers is at least $\frac{1}{414.465}$ .
@@ 1. Introduction
For any positive integers $m$ and $k$, the Waring-Goldbach Problem is to discuss every positive number $\ell_{m,k}$ can be expressed as the sum of the $m$-thth powers of $k$ primes $p_{1},p_{2},\cdots,p_{k}$, i.e.
@@
When $m=1,k=2$, this is the famous Goldbach’s Conjecture, the difficult problem, which has not been successfully solved so far. But when $m=1$, $k=3$, the case of three prime numbers has been proven, which is the famous theorem of three prime numbers[9]. Furthermore, for higher-order terms $m\geq 1$, scholars hope to obtain the smallest $k$ that enables all positive integers to be expressed in the form of (1.1), and denote this smallest $k$ as $G(m)$.
In 1770, Waring deduced from a finite set of evidence that every positive integer is the sum of four squares, nine cubes, nineteen fourth powers, and so on. In 1770, Lagrange proved the existence of $G(2)$, and in the following 139 years, proofs of existence were obtained for $m=3,4,5,6,7,8,10$. In 1909, Hilbert used induction to prove the existence of G(m)G(m)for mm[13].
In addition, for those $m,k$ that cannot express all positive integers in the form of (1.1), scholars also hope to study the density of positive integers that can be expressed in the form of (1.1) among all positive integers in this case. For example $m=3,k=4$, some scholars have already obtained some results. Let $\mathcal{L}_{3,4}$ be the set of integers $\ell_{m,k}$ that can be written as the formula (1.1) when $m=3,k=4$ . In 1949, Roth[5] showed that
@@
Furthermore, from 2001 to 2003, Ren[2][3]improved Roth’s theorem to the extent that
@@
The authors are interested in this problem especially for more cases $m$ and $k$. In this paper we consider the case $m=4,k=7$ and try to say something about it. Let $\ell$ be a positive integer, suppose $\ell$ can be expressed as the sum of the fourth powers of eight prime numbers, i.e.
(1.2) $\ell=p_{1}^{4}+p_{2}^{4}+p_{3}^{4}+\cdots+p_{8}^{4},$
@@
Theorem. $\beta=\frac{1}{414.465}$ is acceptable in (1.3).
Note. Compared to the sums of four cubes of primes, the density of sums of eight fourth power of primes is smaller. For cases with a relatively small number, the results obtained by the method in this paper are not ideal, but if we expand the expression to the sums of more than eight prime numbers, the density may become even higher, which is a problem worthy of further research.
@@ 2. Prerequisite Knowledge
Let $N$ be a large integer, $\delta_{0}$ be small enough and
$U=\left(\frac{N}{64(1+\delta_{0})}\right)^{1/4},\quad V=U^{7/8}.$
@@
(2.1) $J(n)=\int_{-U^{-4}L^{B}}^{U^{-4}L^{B}}\Phi(\lambda,U)\overline{\Psi(\lambda,U)}\Big|\Psi(\lambda,U)\Big|^{6}\Big|\Psi(\lambda,V)\Big|^{8}e(-n\lambda)\>d\lambda.$
@@
(2.3) $T_{d}(n,q)=\sum_{\begin{subarray}{c}a=1\\ (a,q)=1\end{subarray}}^{q}\frac{S(q,ad^{4})C^{7}(q,a)\overline{C^{8}(q,a)}}{q\varphi^{15}(q)}e\biggl(-\frac{an}{q}\biggr)$
@@
(2.4) $S(q,a)=\sum_{m=1}^{q}e\biggl(\frac{am^{4}}{q}\biggr),\quad C(q,a)=\sum_{\begin{subarray}{c}q=1\\ (q,m)=1\end{subarray}}^{q}e\biggl(\frac{am^{4}}{q}\biggr).$
@@ 3. Some Lemmas
Lemma 1. The number S of solutions of
(3.1) $x_{1}^{4}+y_{1}^{4}+y_{2}^{4}=x_{2}^{4}+y_{3}^{4}+y_{4}^{4}$
with $U<x\leq 2U,U^{7/8}<y\leq 2U^{7/8}$ satisfies
$S\ll U^{25/8}.$
Proof. For $x_{1}=x_{2}$, we can easily calculate that the number $S$ of solutions of (3.1) satisfies
$S\ll U^{25/8}.$
For $x_{1}\neq x_{2}$, base on symmetry, we assume $x_{1}<x_{2}$ and write $x_{2}=x_{1}+h$. Then (3.1) becomes
$h(4x_{1}^{3}+6x_{1}^{2}h+4x_{1}h^{2}+h^{3})=y_{1}^{4}+y_{2}^{4}-y_{3}^{4}+y_{4}^{4}.$
Since $y_{1}^{4}+y_{2}^{4}\leq 32U^{7/2}$ and $x_{1}^{3}>U^{3}$ it follows that $h<8U^{1/2}$.
Let
$F_{h}(\alpha)\>=\>\sum_{U<x\>\leqslant\>2U}e\bigl(\alpha h(4x_{1}^{3}+6x_{1}^{2}h+4x_{1}h^{2}+h^{3})\bigr)\>,$
$G(\alpha)=\sum_{h<8U^{1/2}}F_{h}(\alpha)\>,$
$f(\alpha)\>=\>\sum_{U^{7/8}<y\>\leqslant\>2U^{7/8}}e(\alpha y^{4})\>.$
Then
$S\leqslant\int_{0}^{1}G(\alpha)|f(\alpha)|^{4}d\alpha.$
We hope to obtain an upper bound for $G(\alpha)$. Suppose that $|\alpha-a/q|<q^{-2}$ and $(a,q)=1$. By Cauchy’ s inequality
$|G(\alpha)|^{2}\>\ll U^{1/2}\>\sum_{h\><\>8U^{1/2}}|F_{h}(\alpha)|^{2}.$
Moreover, write $y=x+h_{1}$ and we have
$|F_{h}(\alpha)|^{2}\>$ $=\sum_{U<x\>\leqslant\>2U}\sum_{U<y\>\leqslant\>2U}e\bigl(\alpha h(4x_{1}^{3}+6x_{1}^{2}h+4x_{1}h^{2}-4y_{1}^{3}-6y_{1}^{2}h-4y_{1}h^{2}\bigr)$
$=\>\sum_{|h_{1}|<U}\sum_{\max(U,U-h_{1})<x\leqslant\>\min(2U,2U-h_{1})}e\bigl(\alpha h(12h_{1}x^{2}+12h_{1}(h+h_{1})x+4h_{1}^{3}+6hh_{1}^{2}+4h^{2}h_{1})\bigr).$
Record the innermost sum formula as $T(h_{1})$. Use Cauchy’ s inequality again
$|F_{h}(\alpha)|^{2}\>\leq U^{1/2}\biggl(\sum_{|h_{1}|<U}|T(h_{1})|^{2}\biggl)^{1/2}.$
Similarly, write $y^{\prime}=x^{\prime}+h_{1}$, then
$|T(h_{1})|^{2}$ $=\sum_{|h_{2}|<U}\sum_{\max(U,U-h_{2})<x^{\prime}\leqslant\>\min(2U,2U-h_{2})}e(\alpha hh_{1}h_{2}(24x+12(h+h_{1}+h_{2})))$
[…]
Hence
$|F_{h}(\alpha)|^{2}$ $\ll U^{1/2}\biggl(\sum_{|h_{1}|<U}\bigg(U+\sum_{0<h_{2}<U}\min(U,||24\alpha hh_{1}h_{2}||^{-1})\bigg)\biggl)^{1/2}$
[…]
By Lemma 2.2 of Vaughan[7], we have
(3.2) $|G(\alpha)|^{2}\ll U^{5/2+\epsilon}+U^{13/4+\epsilon}q^{-1/2}+U^{11/4+\epsilon}+U^{3/2+\epsilon}q^{1/2}$
Let $\mathfrak{M}^{\prime}(q,a)$ denote the interval $a/q-q^{-1}U^{-2},a/q+q^{-1}U^{-2}]$ and $\mathcal{U}=(P^{-2},1+P^{-2}]$. We may suppose that $U\geqslant 4$. Then the $\mathfrak{M}^{\prime}(q,a)$ with $1\leqslant a\leqslant q\leqslant U$, $(a,q)=1$ are disjoint and contained in $\mathcal{U}$. Let $\mathfrak{M}$ be the union of the $\mathfrak{M}^{\prime}(q,a)$ with $1\leqslant a\leqslant q\leqslant U$, $(a,q)=1$, and let $m^{\prime}=\mathcal{U}\backslash\mathfrak{M}^{\prime}$. Then
$S\leqslant\>\int_{{}_{\mathfrak{M}^{\prime}}}G(\alpha)|f(\alpha)|^{4}d\alpha\>+\>\int_{{}_{m^{\prime}}}G(\alpha)|f(\alpha)|^{4}\>d\alpha.$
By (3.2), when $q>U$, there is $\alpha\notin\mathfrak{M}^{\prime}$ and
$G(\alpha)\ll P^{11/8+\epsilon}.$
Hence, by Lemma 2.5 of Vaughan[7],
$\int_{{}_{m^{\prime}}}G(\alpha)|f(\alpha)|^{4}\>d\alpha\ll P^{25/8}.$
For $\alpha\in\mathfrak{M}^{\prime}$, there is $1\leqslant a\leqslant q\leqslant U$, then by Lemma 6.3 of Vaughan[7]
$G(\alpha)\ll U^{13/8}q^{-1/4},$
$f(\alpha)\ll U^{7/8}q^{-1/4}(1+U^{7/2}|\alpha-a/q|)^{-1}.$
Therefore
$\int_{{}_{\mathfrak{M}^{\prime}}}G(\alpha)|f(\alpha)|^{4}d\alpha$ $\ll\sum_{q\leq U}\sum_{\stackrel{{\scriptstyle a=1}}{{(a,q)=1}}}^{q}U^{13/8+7/2}q^{-1-1/4}\int_{\mathfrak{M}^{\prime}(q,a)}\left(1+U^{7/2}\bigg|\alpha-\frac{a}{q}\bigg|\right)^{-4}d\alpha$
$\ll U^{19/8}$
This proves Lemma 1.
@@
$p_{2},p_{3},p_{4},p_{9},p_{10},p_{11},p_{12}\sim U,\quad p_{5},p_{6},p_{7},p_{8},p_{13},p_{14},p_{15},p_{16}\sim V.$
For $0<\xi<9/25$ and $D=N^{\xi}$, take $\mathfrak{S}_{d}(n)$ and $J(n)$ from the above (2.1), (2.2), define $E_{d}(n)$ as follows:
@@
$\ll\int_{U^{-4}L^{B}}^{\infty}|\Phi(\lambda,U)||\Psi(\lambda,U)|^{7}|\Psi(\lambda,V)|^{8}\>d\lambda$
$\ll U^{-24}V^{-24}L^{-15}\int_{U^{-4}L^{B}}^{\infty}\min(U^{4},|\lambda|^{-1})^{6}\min(V^{4},|\lambda|^{-1})^{8}\>d\lambda$
$\ll U^{4}V^{8}L^{-5B}.$
@@
$\int_{-\infty}^{\infty}\Phi(\lambda,U)\overline{\Psi(\lambda,U)}\big|\Psi(\lambda,U)\big|^{6}\big|\Psi(\lambda,V)\big|^{8}e(-n\lambda)\>d\lambda$
$=$ $\frac{1}{4}\int_{\mathcal{D}}\frac{d\nu_{1}\cdots d\nu_{8}du_{1}\cdots du_{7}}{\nu_{1}^{3/4}\cdots\nu_{8}^{3/4}u_{1}^{3/4}\cdots u_{8}^{3/4}\log\nu_{1}\cdots\log\nu_{8}\log u_{1}\cdots\log u_{7}},$
@@
$\mathcal{D}=\left\{(v_{1},\ldots,v_{8},u_{1},\ldots,u_{7})\colon V^{4}\leqslant v_{1},\ldots,v_{8}\leqslant 16V^{4},\ U^{4}\leqslant u_{1},\ldots,u_{7}\leqslant 16U^{4}\right\}$
and $u_{8}=n+v_{1}+v_{2}+v_{3}+v_{4}-v_{5}-v_{6}-v_{7}-v_{8}+u_{1}+u_{2}+u_{3}+u_{4}-u_{5}-u_{6}-u_{7}.$
@@
(3.3) $\begin{split}J(n)=&\int_{-\infty}^{\infty}\Phi(\lambda,U)\overline{\Psi(\lambda,U)}|\Psi(\lambda,U)|^{6}|\Psi(\lambda,V)|^{8}e(-n\lambda)d\lambda\\ &+\>O(U^{4}V^{8}L^{-5B})\\ =&\frac{1}{4}\int_{V^{4}}^{16V^{4}}\>\frac{dv_{1}}{v_{1}^{3/4}\log v_{1}}\cdots\int_{V^{4}}^{16V^{4}}\>\frac{dv_{4}}{v_{8}^{3/4}\log v_{8}}\int_{U^{4}}^{16U^{4}}\>\frac{du_{1}}{u_{1}^{3/4}\log u_{1}}\\ &\times\int_{U^{4}}^{16U^{4}}\frac{du_{2}}{u_{2}^{3/4}\log u_{2}}\cdots\int_{\max(U^{4},x-16U^{4})}^{\min(16U^{4},x-U^{4})}\frac{du_{7}}{u_{7}^{3/4}(x-u_{7})^{3/4}\log u_{7}}\\ &+\>O(U^{4}V^{8}L^{-5B}).\end{split}$
Let the last integral in the equation be denoted as $I$, where $x=n+v_{1}+v_{2}+v_{3}+v_{4}-v_{5}-v_{6}-v_{7}-v_{8}+u_{1}+u_{2}+u_{3}+u_{4}-u_{5}-u_{6}$, now we calculate the upper bound of $I$. Firstly, regarding $x\leqslant 2U^{4}$ or $x\geqslant 32U^{4}$, in both cases, the integral region does not exist, so we have $2U^{4}<x<32U^{4}$. In this case, we record as $u_{7}=xu$, then there is